OpenStudy (anonymous):

William needs a loan for $7550. He can borrow that amount at 4% interest for a 3-year term or at 5% interest for a 3-year term. How much more interest will William pay if he borrows the money at 5% for a 3-year term?

OpenStudy (kash_thesmartguy):

\[I=prt\]\[(I=0.05*7500*3)-(I=0.04*7500*3)\]\[1125-900=225\] William will pay $225 more if he chooses to borrow the money at 5% for a 3-year term.